OpenStudy (anonymous):

Which has a greater area, four circles, each with radius 1 m, or one circle with radius 4 m? Area= Pir2 so are the circles the same or is the one circle greater? kinda confused dont know which one it is

Directrix (directrix):

If you use the area of a circle formula and take a look at the area of a circle of radius 4, and then calculate the area of a circle of radius 1 and quadruple the result of that, it will be clear as to which is larger.

Directrix (directrix):

Compare A = pi * (4)^2 with A = 4 * pi * (1)^2. Are the results the same? @sharlener

OpenStudy (anonymous):

@Viyang r u solving the question or .............................. what lolololol

OpenStudy (anonymous):

No

OpenStudy (anonymous):

the larger circle is greater

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

wow viyang!! i know ur intentions

Directrix (directrix):

@sharlener >>the larger circle is greater Yes, the radius 4 is squared.